OpenStudy (anonymous):

evaluate the integral (dy/(4-y^2))

OpenStudy (lalaly):

\[\frac{1}{4-y^2} = \frac{1}{(2-y)(2+y)}\] now break into partial fractions

OpenStudy (amistre64):

i was kinda partial to it too

OpenStudy (lalaly):

\[\frac{1}{4-y^2} = \frac{A}{2-y}+\frac{B}{2+y}\]\[1=A(2+y)+B(2-y)\]\[1=2A+Ay+2B-By\]\[2A+2B=1\]\[A-B=0\]\[A=B\]\[4A=1\]\[A=B=\frac{1}{4}\] \[\int\limits{\frac{dy}{4-y^2}}=\int\limits{\frac{dy}{4(2-y)}}+\int\limits{\frac{dy}{4(2+y)}}\]

OpenStudy (lalaly):

can u integrate now?

OpenStudy (anonymous):

well couldnt we just use substitution

OpenStudy (amistre64):

lol .... and sub into what? y^2 doesnt derive into a constant

OpenStudy (anonymous):

no i mean using u substitution ?

OpenStudy (anonymous):

partial fractions my favorite! you know how to do it the french way?

OpenStudy (lalaly):

\[\frac{1}{4} \int\limits{\frac{dy}{2-y}}+\frac{1}{4}\int\limits{\frac{dy}{2+y}} = \frac{1}{4} (\ln(2-y)+\ln(2+y))\]

OpenStudy (anonymous):

you want \[\frac{1}{(2-x)(2+x)}=\frac{A}{2-x}+\frac{B}{2+x}\] so find A think " x cannot be 2} to take \[\frac{1}{(2-x)(2+x)}\] get rid of the factor x-2 and put \[\frac{1}{2+x}\] replace x by 2 and get \[A=\frac{1}{4}\]

OpenStudy (anonymous):

you can do it in your head by putting your finger over the factor you don't want and substituting in the number. french guy showed me that easy trick

OpenStudy (lalaly):

i like it :D